OpenStudy (anonymous):

how can i solve these with substitution? i'm pretty good at solving these with the elimination method, but i need practice using substitution. x+4y=0 3x+2y=20

OpenStudy (anonymous):

Since you have two sets of functions, you can substitute another by partially solving one (such as solve x+4y=0, as far as possible), and applying it to the other and solve for y. After doing that, solve for x.

OpenStudy (anonymous):

solve the first equation for x, so x = -4y now plug this into the second equation: you will get: 3(-4y) + 2y = 20

OpenStudy (anonymous):

I just typed a response and realized my math was off... but here is the correct way. Solve the first equation for x. You always want to get one of the equations down to one variable. It doesn't matter which equation. In this one the first equation is easiest. x=-4y Plug the value of x into the second equation: 3(-4y)+2y=20 -12y+2y=20 -10y=20 y=-2 And then, plug the y value into the first equation to solve for x. x+4(-2)=0 x-8=0 x=8 (8,-2)

OpenStudy (anonymous):

ohhh okay thank you everyone! <3 those negatives are what keep throwing me off in that second equation, i keep putting x+4-2=0 instead of x+4(-2)=0